A 0.11 kg bullet traveling at speed hits a 18.3 kg block of wood and stays in the wood. The block with the bullet imbedded in it moves forward with a velocity of 8.8 m/s. What was the velocity (speed) of the bullet immediately before it hit the block (in m/s)?
Answer #1

using conservation of momentum

m1 v1 = (m1 + m2) v

0.11* v1 = (18.3 + 0.11)* 8.8

v1 = 1472.8 m/s

======

Comment before rate in case any doubt, will reply for sure.. goodluck